# On the dynamics of leapfrogging vortex rings

### v1

@@ Abstract.
The evolution of highly concentrated vorticity around rings in the three-dimensional axisymmetric Euler equations is studied in a regime for which the leapfrogging dynamics predicted by Helmholtz is expected to occur. For a general class of initial data corresponding to vortex rings of sufficiently small thickness $\varepsilon>0$, we prove weak and strong localization estimates for the vortex rings at positive times up to some time $T_{0}/|\ln\varepsilon|$ for fixed $T_{0}$, independent of $\varepsilon$. Moreover, we derive the motion law for the position of the rings on $[0,T_{0}/|\ln\varepsilon|]$.
To the best of our knowledge, this is the first result of a rigorous derivation of the leapfrogging motion up to times of order $|\ln\varepsilon|^{-1}$ for general initial data and the suitable scaling regime. The singular interaction of rings requires significant improvements of weak and strong localization estimates obtained in prior works. Our method is based on the combination of a new variational argument and a recently introduced double iterative procedure. Possibly due to natural filamentation of the vortex rings, which induces a lack of control of the support of the solution in the symmetry axis direction, our method does not allow to reach the time at which the rings have performed a full rotation around each other.
@@ 1. Introduction
As a model of the evolution of an incompressible and inviscid three-dimensional fluid, we consider the standard 3D Euler equations
where $\mathbf{U}$ is uniquely determined by $\mathbf{\Omega}$ for all $t\in[0,T]$ by means of the Biot-Savart law (the second equation in (1.2)). In the present paper, we restrict our attention to solutions that are axisymmetric vector fields without swirl, namely if considered in cylindrical coordinates $(e_{z},e_{r},e_{\theta})$, the solution $\mathbf{U}$ satisfies
$\mathbf{U}(z,r,\theta)=U_{z}(z,r)e_{z}+U_{r}(z,r)e_{r},$
$\mathbf{\Omega}=\Omega e_{\theta},\quad\text{with}\,\,\Omega(z,r)=\operatorname{curl}\begin{pmatrix}U_{z}\\ U_{r}\end{pmatrix}=(\partial_{z}U_{r}-\partial_{r}U_{z})(z,r).$
Problem (1.2) can hence be reduced to a two-dimensional problem for the potential scalar vorticity $\omega=r^{-1}\Omega$ posed on the half plane $(z,r)\in{\mathbb{R}}_{+}^{2}={\mathbb{R}}\times{\mathbb{R}}_{+}$, see Section 2 for details:
(1.3) $\begin{cases}\partial_{t}\,\omega+u\cdot\nabla\omega=0&\text{ in }{\mathbb{R}}_{+}^{*}\times{\mathbb{R}}^{2}_{+},\vspace{1mm}\\ \operatorname{div}(ru)=0&\text{ in }{\mathbb{R}}_{+}\times{\mathbb{R}}^{2}_{+},\vspace{1mm}\\ \operatorname{curl}(u)=r\omega&\text{ in }{\mathbb{R}}_{+}\times{\mathbb{R}}^{2}_{+},\vspace{1mm}\\ \omega(0,\cdot)=\omega_{0}&\text{ in }{\mathbb{R}}^{2}_{+}.\end{cases}$
The scope of this paper is to investigate the dynamics when the vorticity is sharply concentrated around rings. Specifically, we consider initially concentrated vortex rings of thickness $\varepsilon$, for sufficiently small $\varepsilon$ and we rigorously identify the motion of rings at leading order. The study of such vortex rings goes back to Helmholtz [29] and J. J. Thomson [48]. This is the subject of many studies ranging from mathematics to engineering. They are observed in 3D flows with high Reynolds number as long-lived vortex structures that translate along their symmetry axis. It was computed by Kelvin [36] that a vortex ring of major radius $r$, minor radius $\varepsilon$, and circulation $\gamma$ has a self-induced velocity
(1.4) $v=\frac{\gamma}{4\pi r}|\ln\varepsilon|e_{z},$
at the first order in $\varepsilon$, see also [25, 5] for rigorous derivation. The respective motion law for a single vortex ring was rigorously established for weakly viscous fluids in [7, 9, 27]. More generally, assuming that the vorticity is concentrated around a curve, the motion is expected to be asymptotically governed by the binormal curvature flow equation, as was formally computed by Da Rios [47]. The binormal curvature flow exhibits a rich panel of dynamics that have been intensively studied. In particular a serie of papers by Banica and Vega is devoted to self-similar solutions or infinite polygonal lines solutions (we refer to [4] for a review of the results). Finally, Jerrard and Smets [34] introduced a new formulation of the binormal curvature flow equation allowing for weaker solutions.
The issue of the persistence in time of such vortex filaments by the Euler flow is the so-called vortex filament conjecture, see [33, 24] for partial results in this direction. While the conjecture remains open in its generality, results are available for specific geometric and dimension reduced settings. The rigorous asymptotic dynamics of 2D point vortices (the motion of Dirac masses of vorticity in the 2D Euler equations) was obtained in [40]. In the aforementioned case of annular vortex rings evolving according to (1.3), positive answers have been given with various assumptions. We refer to the work by Buttà, Cavallaro and Marchioro [8, 11] and references therein. This is in particular due to the fact that the problem allows for a dimension reduction leading to the 2D problem (1.3).
We emphasize that the system (1.3) notably differs from the 2D Euler equations. Indeed, the dynamics of 2D Euler point-vortices do not include any self-induced motion of vortices. Even though the problem (1.3) is reduced to two dimensions, as in the study of helical filaments in [20] or point-vortices for the lake equations [31], the Biot-Savart law (the relation giving $u$ in terms of $\omega$) is much more complex compared to the planar case. Additional techniques are required and the results developed for the 2D Euler equations (see for instance [40]) cannot be applied. Roughly speaking, the motion of vortex rings is driven by the self-induced motion (1.4) and by their mutual interactions. Depending on the chosen asymptotic regime, the former maybe dominant, leading to a translation of the vortex rings along the symmetry axis [12, 8]. In case the latter dominates, the dynamics is reminiscent to the 2D dynamics of point vortices [10]. In particular, it was shown in [8] that the interactions of vortex rings of different radii (independent of $\varepsilon$) are bounded independently of $\varepsilon$ and can be neglected compared to their self-interaction (inducing the velocity given by (1.4)) which is of order $|\ln\varepsilon|$.
However, it is well-known since the work of Helmholtz [29], see also [21, 22, 30, 37, 6], and it has been confirmed by numerical [3, 38, 45] and experimental evidence [51] that in particular conditions, two coaxial vortex rings can successively pass through each other in a periodic motion known as the leapfrogging phenomenon. For such phenomena to occur provided, the vortex rings need to be located at infinitesimal mutual distance leading to mutual interactions inducing a change of their radii that is significant enough to cause a variation of their self-induced velocity (1.4). We say that leapfrogging dynamics occurs when the variation of vertical velocity in the self-interaction is of the same order as the mutual interactions of ring. It turns out that this happens when the vortex rings are placed a distance of order $1/\sqrt{|\ln\varepsilon|}$ to each other, and a detailed explanation of the suitable scaling regime is provided in Appendix B, see also [18, 35]. In the recent paper [18], the authors construct a smooth solution to (1.3) exhibiting the leapfrogging motion. The 3D Gross-Pitaevskii equation also exhibits leapfrogging dynamics as proven by Jerrard and Smets [35].
In the present paper, we rigorously derive the leapfrogging dynamics for (1.3) in the aforementioned scaling regime and a general class of initial data that remained open despite being expected since [29], see Theorem 1.1 below. We point out that the present scaling regime differs from the one in [11] in which the vortex positions are located at a mutual distance of order $1/|\ln\varepsilon|$. While the mutual interactions are more singular compared to our setting, the estimates obtained there hold on a significantly shorter time interval. The differences between the regimes are discussed in detail in Appendix B.
In the present setting, all vortex rings move at leading order at the same vertical velocity $v\propto|\ln\varepsilon|$ (relation (1.4)), at a relative distance of order $1/\sqrt{|\ln\varepsilon|}$ and in a relative motion of velocity $v^{\prime}\propto\sqrt{|\ln\varepsilon|}$. Therefore, it is very natural to rescale the time variable by a factor $|\ln\varepsilon|$, so that both motions happen on a time of order 1. More precisely, we introduce the system
@@
(1.5) $\begin{cases}\displaystyle\partial_{t}\,\omega^{\varepsilon}+\frac{1}{|\ln\varepsilon|}u^{\varepsilon}\cdot\nabla\omega^{\varepsilon}=0&\text{ in }{\mathbb{R}}_{+}^{*}\times{\mathbb{R}}^{2}_{+},\vspace{1mm}\\ \operatorname{div}u^{\varepsilon}=0&\text{ in }{\mathbb{R}}_{+}\times{\mathbb{R}}^{2}_{+},\vspace{1mm}\\ \operatorname{curl}(u^{\varepsilon})=r\omega^{\varepsilon}&\text{ in }{\mathbb{R}}_{+}\times{\mathbb{R}}^{2}_{+},\vspace{1mm}\\ \omega^{\varepsilon}(0,\cdot)=\omega_{0}^{\varepsilon}&\text{ in }{\mathbb{R}}^{2}_{+}.\end{cases}$
for which we expect two vortex rings at a distance of order $1/\sqrt{|\ln\varepsilon|}$ would perform a complete rotation around each other in a time of order 1, while traveling altogether a vertical distance also of order 1.
@@
We can now formulate our problem in a precise mathematical framework and state our result. For $x\in{\mathbb{R}}^{2}_{+}$, we further denote by $x=(x_{z},x_{r})$ with $x_{z}\in{\mathbb{R}}$ and $x_{r}\in{\mathbb{R}}_{+}$, and by $(e_{z},e_{r})$ the canonical basis on ${\mathbb{R}}^{2}$. We consider a fixed reference point:
@@
For $\varepsilon>0$, the initial data $\omega_{0}^{\varepsilon}$ will be a superposition of $N$ vorticities sharply concentrated around points located at distance of order $1/\sqrt{|\ln\varepsilon|}$ from the reference point:
@@
where $Y_{i,0}\in{\mathbb{R}}^{2}$ is fixed independently of $\varepsilon$. We next assume that the vortex filaments all have the same density $\gamma$. Namely, let $\gamma\neq 0$, we assume that $\omega_{0}^{\varepsilon}\in L^{\infty}({\mathbb{R}}^{2}_{+})$ such that
@@
for some $M_{0}>0$ independent of $\varepsilon$. The goal of this paper is to prove that the sequence of vorticities $\omega^{\varepsilon}(t,\cdot)$, solution of (1.5) for such intial data, remains asymptotically concentrated around $N$ points $X_{i}^{\varepsilon}(t)$ and that the trajectories of these points is governed by the ODE:
Before stating our result rigorously, we briefly comment on the ODE (1.8). We recovered, in the first term, the vertical self-induced velocity (1.4), and in the second term, the well-known mutual interaction in the 2D point-vortex dynamics, both up to a factor $1/|\ln\varepsilon|$ coming from the time rescaling. The system (1.8) is clearly Hamiltonian. It is not known whether the solution is global, should several vortex rings collapse together at $r=0$. In Section 5.2, we prove the existence of a time $T_{X}$ depending on $(Y_{i,0})_{1\leqslant i\leqslant N}$ but not on $\varepsilon$ such that the system (1.8) has a solution on the time interval $[0,T_{X}]$. To this end, we observe that $t\mapsto Y_{i}^{\varepsilon}(t)$ defined by the relation
(1.9) $X_{i}^{\varepsilon}(t)=X^{*}+\frac{\gamma}{4\pi r^{*}}te_{z}+\frac{1}{\sqrt{|\ln\varepsilon|}}Y_{i}^{\varepsilon}(t),$
converge as $\varepsilon\to 0$ to a limiting trajectory $t\mapsto\widetilde{Y_{i}}(t)$. This limiting trajectory satisfies
[…]
and in works such as [35, 18, 42], it is this dynamics that is considered as the motion of the rings in the right scaling. For (1.10), no collisions occur in finite time, see Appendix C. In this paper we prefer to work with the trajectories $(X_{i})_{1\leqslant i\leqslant N}$ and $(Y_{i})_{1\leqslant i\leqslant N}$ as it is much more natural in the computations, then prove in Section 5 that in all the following, one could replace the trajectories $t\mapsto X_{i}^{\varepsilon}(t)$ by
to match the notations of [35, 18, 42], and all the results, in particular the following Theorem 1.1, would still hold.
We are now in position to give precisely our main theorem.
@@
There exists $T_{0}>0$, $\varepsilon_{0}>0$ and $C_{0}>0$, depending only on $X^{*}$, $(Y_{i,0})$ and $M_{0}$, such that for every $\varepsilon\in(0,\varepsilon_{0}]$, there exists a unique solution $(t\mapsto X_{i}^{\varepsilon})_{1\leqslant i\leqslant N}$ of (1.8) on $[0,T_{0}]$, and the unique weak solution $(u^{\varepsilon},\omega^{\varepsilon})$ of (1.5) on $[0,T_{0}]$ (in the sense of Proposition  2.1 below) can be decomposed as
@@
with the following properties:
@@
Let us now formulate several important remarks regarding Theorem 1.1. Concerning previous results and to the best of our knowledge, the only rigorous results concerning the leapfrogging dynamics for the 3D Euler equations are [18, 11] and [35] for the 3D Gross-Pitaevskii equation. While [18, 35] consider the same scaling regime, the regime studied in [11] differs from the present one, as previously mentioned. The results from [11] cannot be adapted to our situation, as also pointed out by the authors in [11, p.791]. In [18], the authors construct initial data $\omega_{0}^{\varepsilon}$ of a very specific class to obtain a smooth solution exhibiting leapfrogging vortex rings. We wish to emphasize the fact that the conclusions of Theorem 1.1 hold for any initial data $\omega_{0}^{\varepsilon}\in L^{\infty}({\mathbb{R}}^{2}_{+})$ satisfying relations (1.7), which ranges from smooth compactly supported blobs to vortex patches, even possibly multiply connected. Opposite to [18], our proof does not rely on sophisticated elliptic desingularization techniques. The method in [35] is based on Jacobian estimates not available for (1.1).
The weak localization property, namely item $(i)$ in Theorem 1.1, is obtained by a new and original argument. Assuming weak and strong localization, we obtain a precise expansion of the vortex rings’ energy from which we deduce a concentration property, inspired by the work of Turkington in [49]. By doing so, we largely improve the previously known weak localization estimates [8]. Proposition 4.1 below provides even more precise estimates than what is stated in Theorem 1.1. In particular, for any $\eta>0$ it gives the existence of a domain of size $\varepsilon^{1-\eta}$ containing most of the potential vorticity. However, interestingly enough, our method fails to provide a precise location for this domain, see Remark 4.13 for a detailed discussion. Finally, a bootstrap argument yields that the weak concentration holds as long as the strong localization does.
The strong localization property $(ii)$ in Theorem 1.1 is pivotal for deriving the asymptotic motion and constitutes a substantial improvement compared to previously known results in [8] and similarly in [31, 20] in different geometric settings. These previous results state that the potential vorticity remains supported in a strip of radial size $1/|\ln\varepsilon|^{\kappa}$, for $\kappa<1/4$. Here, we prove a localization of radial size $1/\sqrt{|\ln\varepsilon|}$. This constitutes the critical size in order to observe a motion on spatial scales of order $1/\sqrt{|\ln\varepsilon|}$ on which leapfrogging occurs. This is achieved in two independent steps. We first prove the confinement in the radial direction by adapting the double iterative procedure recently introduced in [11] to the present regime. Second, we prove the localization in the vertical direction in a strip of size of order $1$ being less precise than the radial direction and therefore easier to obtain. On a related note, in the case of a single vortex ring, the result of Theorem 1.1 also improves those of [8].
In the following Remark 1.2, the localization property and its link to the small-time assumption for our main result are discussed.
@@
A natural and relevant question regarding our main result Theorem 1.1 is whether the small time assumption can be removed as it prevents to describe a full rotation of the vortex rings around each other. In [18], a particular smooth solution to (1.3) exhibiting the leapfrogging dynamics for an arbitrarily large number of rotations is constructed. In [11], an arbitrarily large number of rotations of two vortex rings around each other is observed provided that the vortex rings radii’ are large enough. As detailed in Appendix B, translating the result of [11] to the present setting by means of the scaling invariance of (1.3) yields a time scale of order $|\ln\varepsilon|^{-2}$ being much shorter than the present one.
Concerning possible generalization of our result, we expect that our small time assumption is not sharp and can be optimized. However, it becomes clear from our proof that our method is not expected to reach a full rotation. We only obtain a $\mathcal{O}(1)$ localization in the $z$ direction, and the strips in which each ring is supported are not allowed to overlap. Moreover, additional assumptions on the general class of initial data (1.7) considered here might be necessary. Indeed, it is conceivable that the possible occurrence of filamentation of the vortex ring in the vertical direction prevents an improvement of the $\mathcal{O}(1)$ localization in the $z$ direction for general initial data. Namely, a tail in the wake of the ring may appear, a phenomenon that was described for instance in [43, 44] and recently proved in [16] for Hill’s vortex. This fact may explain similar restrictions in [8, 12, 20, 31]. In [18], this filamentation is not present since the authors work with specific initial data and a stable vorticity profile is constructed. In [11], this filamentation does also not occur suggesting that the tail does appear on the short time scale considered.
Nevertheless, it might be possible to remove the small time assumption at least for the weak localization: this would require the weak localization to be obtained without assuming a priori strong localization, which appears not to be withing reach of our current method. For these reasons, the question of the stability of a concentrated vortex when colliding with the tail of another vortex seems to be an important and challenging direction for future research, which will require new tools.
An interesting quantity to estimate is the center of potential vorticity of each vortex blob in the 2D reduction, namely the expression
(1.12) $b^{\varepsilon}_{i}(t)=\int_{{\mathbb{R}}^{2}_{+}}x\frac{x_{r}\omega^{\varepsilon}_{i}(t,x)}{\gamma}\mathrm{d}x.$
We obtain in Proposition 5.4 as a consequence of the localization properties that

### v2

@@ Abstract.
The evolution of highly concentrated vorticity around rings in the three-dimensional axisymmetric Euler equations is studied in a regime for which the leapfrogging dynamics predicted by Helmholtz is expected to occur. We provide in this paper the first result deriving this phenomenon for a general class of initial data in the suitable regime. The singular interaction of rings requires significant improvements of weak and strong localization estimates obtained in prior works. Our method is based on the combination of a new variational argument and a recently introduced double iterative procedure.
@@ 1. Introduction
The goal of this paper is to rigorously study the evolution of interacting vortex rings in $3D$ incompressible ideal fluids. Vortex rings are experimentally and numerically well observed coherent structures propagating in an ideal fluid. Mathematically, they correspond to axisymmetric solutions to the 3D incompressible Euler equations concentrated around annuli. Due to its self-induced motion, linked to local induction approximation (LIA) [54], a single vortex ring translates at leading order along its symmetry axis. Specifically, it was computed by Kelvin [40, 41] that a vortex ring of major radius $r$, thickness $\varepsilon$, and circulation $\gamma$ has a self-induced velocity at leading order equal to
(1.1) $v=\frac{\gamma}{4\pi r}|\ln\varepsilon|e_{z}.$
It is well-known since the work of Helmholtz [33] in 1858, see also [25, 26, 34, 41, 6], that two coaxial vortex rings can successively pass through each other in a periodic motion. More precisely and according to [33], two given vortex rings translating along the same axis may exhibit the so-called leapfrogging motion in which ‘the foremost widens and travels more slowly, the pursuer shrinks and travels faster till finally, if their velocities are not too different, it overtakes the first and penetrates it. Then the same game goes on in the opposite order, so that the rings pass through each other alternately’, see Figure 1.
$z$$z$$z$$z$
Figure 1. Leapfrogging phenomenon
For such phenomena to occur, suitable assumptions on the initial locations, radii $r$ and thickness $\varepsilon$ and intensity $\gamma$ of the coaxial vortex rings need to be made. Specifically, coaxial vortex rings need to be located at very small mutual distance leading to sufficiently strong interactions inducing a change of their radii at a suitable scale. This in turn causes a relevant variation of their self-induced velocity (1.1). Here, we say that leapfrogging dynamics occurs when the variation of the self-induced motion (1.1) is of the same order as the mutual interaction of the vortex rings, namely $\sqrt{|\ln\varepsilon|}$. We refer the reader to Appendix B for a detailed discussion of the suitable scaling regime and motion law.
Up to a few crossings, such motion can be experimentally [58] and numerically [3, 42, 52] observed, supporting this conjecture for finite times.
The suitable framework to set up the mathematical problem is given by the axisymmetric Euler equations without swirl in vorticity form in cylindrical coordinates $(z,r,\theta)$. After dimension reduction due to the axial symmetry and for the potential vorticity $\omega_{\varepsilon}$, see Section 2 below for details, we are led to the system
@@
(1.2) $\begin{cases}\displaystyle\partial_{t}\,\omega^{\varepsilon}+\frac{1}{|\ln\varepsilon|}u^{\varepsilon}\cdot\nabla\omega^{\varepsilon}=0&\text{ in }{\mathbb{R}}_{+}^{*}\times{\mathbb{R}}^{2}_{+},\vspace{1mm}\\ \operatorname{div}(ru^{\varepsilon})=0&\text{ in }{\mathbb{R}}_{+}\times{\mathbb{R}}^{2}_{+},\vspace{1mm}\\ \operatorname{curl}(u^{\varepsilon})=r\omega^{\varepsilon}&\text{ in }{\mathbb{R}}_{+}\times{\mathbb{R}}^{2}_{+},\vspace{1mm}\\ \omega^{\varepsilon}(0,\cdot)=\omega_{0}^{\varepsilon}&\text{ in }{\mathbb{R}}^{2}_{+},\end{cases}$
where ${\mathbb{R}}_{+}^{2}$ denotes the space $\{(z,r)\in{\mathbb{R}}\times{\mathbb{R}}_{+}\}$.
In the present setting, all vortex rings obey at leading order the same self-induced motion (1.1). This suggests to scale the time variable by a factor $1/|\ln\varepsilon|$ yielding (1.2). The mutual distance of vortex rings is then chosen to be of order $1/\sqrt{|\ln\varepsilon|}$ leading to a relative motion of velocity $v^{\prime}=\mathcal{O}\left(\sqrt{|\ln\varepsilon|}\right)$. Indeed, in a time of order $1$ for this system, the vortex rings are expected to exhibit a complete rotation around each other while traveling a vertical distance of order $1$, see also Appendix B.
In a slightly informal way, the leapfrogging conjecture can be formulated as follows.
Conjecture 1.
Let $N>1$ and $0<\varepsilon\ll 1$. Given initial vorticity $\omega_{0}^{\varepsilon}$ decomposed in $N$ blobs, each of circulation $\gamma>0$ and sharply ‘concentrated’ around $N$ points $X_{i,0}^{\varepsilon}$ (in terms of $\varepsilon$) located at mutual distance of order $1/\sqrt{|\ln\varepsilon|}$, the unique solution $\omega^{\varepsilon}(t,\cdot)$ to (1.2) remains ‘concentrated’ around $N$ points $X_{i}^{\varepsilon}(t)$.
Moreover, the trajectories of these points are governed by the system
The meaning of ‘concentration’ will be precised below. Concerning (1.3), the motion of the vortex centers is driven by the vertical self-induced velocity corresponding to the first term and the second term that represents the well-known mutual interaction in the 2D point-vortex dynamics. Due to the chosen fast time scale, both are reported up to a factor $1/|\ln\varepsilon|$. The choice of placing the vortex centers at mutual distance of order ${1}/{\sqrt{|\ln\varepsilon|}}$ yields that the relative self-induced motion, i.e. the second order term in (1.1), is of the same order than the motion due to the interaction, see Section 1.2. Note that the velocity induced by the mutual interactions is inverse proportional to the inter-vortex distance. In the case $N=2$, System (1.3) describes the evolution of two leapfrogging vortex rings.
The key difficulties in attempting to prove this conjecture are twofold. The present regime is critical in the sense that the aforementioned effects are of the same strength and the vortex centers are situated very close to each other. This makes the problem at hand conceivably harder than the ones in which the motion is dominated by either of these effects. Commonly, showing the persistence of the localization property on suitable time-scales turns out to be the major difficulty. Second, the asymptotic motion law for small $\varepsilon>0$ needs to be derived. Experimental and numerical evidence supports the conjecture up to a few number of crossings. However, in these experiments various instability mechanisms are displayed and the periodic motion does not occur for arbitrarily large times [42, 47, 52, 57, 58]. The vortex filamentation phenomenon appears to be particularly relevant, it causes a vortex ring to leave a small tail of vorticity behind which may potentially collide with following vortex rings, see for this mechanism in various scenarios [15, 16, 24, 49, 51].
The main result of this paper validates the above conjecture: we prove that an arbitrary but finite number $N$ of initially sharply concentrated vortex rings remains concentrated for a small but positive time independent of $\varepsilon>0$. The initial concentration assumptions allow for general data to be considered, and in particular, do not require any specific vorticity profile.
1.1. Main result
@@
We introduce the precise mathematical setup for the problem and state our main result. For $x\in{\mathbb{R}}^{2}_{+}$, we denote by $x=(x_{z},x_{r})$ with $x_{z}\in{\mathbb{R}}$ and $x_{r}\in{\mathbb{R}}_{+}$, and by $(e_{z},e_{r})$ the canonical basis on ${\mathbb{R}}^{2}$. We consider a fixed reference point
@@
For $\varepsilon>0$, the initial data $\omega_{0}^{\varepsilon}$ will be given by a superposition of $N$ vorticities of the same circulation $\gamma$ and sharply concentrated around points located at distance of order $1/\sqrt{|\ln\varepsilon|}$ from the reference point
@@
where $Y_{i,0}\in{\mathbb{R}}^{2}$ is fixed independently of $\varepsilon$. Specifically, let $\gamma\neq 0$ and assume that $\omega_{0}^{\varepsilon}\in L^{\infty}({\mathbb{R}}^{2}_{+})$ such that
@@
for some $M_{0}>0$ independent of $\varepsilon$. Given that the potential vorticity $\omega^{\varepsilon}$ satisfies the transport equation (1.2) and the existence of a unique associated flow, see Remark 2.3, it also holds $\omega^{\varepsilon}(t)=\sum_{i=1}^{N}\omega_{i}^{\varepsilon}(t)$ for $t\geqslant 0$. Further, it follows from the transport equation that $\omega^{\varepsilon}(t)$ conserves the sign definiteness as well as the $L^{\infty}$-norm. By the divergence free condition of $ru^{\varepsilon}$, (1.2) also implies the conservation of the mass of each blob:
+$\int_{{\mathbb{R}}^{2}_{+}}x_{r}\omega_{i}^{\varepsilon}(t,x)\mathrm{d}x=\gamma,\quad\forall t\geqslant 0,$
see Section 2.2 and (3.3) for more details.
We are now in position to precisely formulate our main result.
@@
There exists $T_{0}>0$, $\varepsilon_{0}>0$ and $C_{0}>0$, depending only on $X^{*}$, $(Y_{i,0})_{1\leqslant i\leqslant N}$ and $M_{0}$, such that for every $\varepsilon\in(0,\varepsilon_{0}]$, there exists a unique solution $(t\mapsto X_{i}^{\varepsilon})_{1\leqslant i\leqslant N}$ of (1.3) on $[0,T_{0}]$, and the unique weak solution $(u^{\varepsilon},\omega^{\varepsilon})$ of (1.2) on $[0,T_{0}]$ (in the sense of Proposition  2.1 below) can be decomposed as
@@
with the following properties
@@
We emphasize that the conclusions of Theorem 1.1 hold for any initial data $\omega_{0}^{\varepsilon}\in L^{\infty}({\mathbb{R}}^{2}_{+})$ satisfying assumptions (1.5), which ranges from smooth compactly supported blobs to vortex patches, even possibly multiply connected. No further assumptions on the regularity or geometric properties of the data are required.
The weak localization property, namely item $(i)$ in Theorem 1.1 states that most of vorticity’s mass is contained in a ball centered in $X_{i}^{\varepsilon}$ obeying (1.3) and provides precise rates of convergence. We largely improve the previously known weak localization estimates [8] and similarly in [35, 23] in different geometric settings and references therein. Proposition 4.1 below provides even more precise estimates than Theorem 1.1. Specifically, for any $\eta>0$ it gives the existence of a domain of size $\varepsilon^{1-\eta}$ containing most of the potential vorticity. However, interestingly enough, our method fails to provide a precise location for this domain, see Remark 4.13 for a detailed discussion.
The strong localization property $(ii)$ in Theorem 1.1 is pivotal for deriving the asymptotic motion and constitutes a substantial improvement compared to previously known results in [8, 23, 35]. These previous results state that the potential vorticity remains supported in a strip of radial size $1/|\ln\varepsilon|^{\kappa}$, for $\kappa<1/4$. Here, we prove a localization of radial size $1/\sqrt{|\ln\varepsilon|}$. This improvement is crucial to reach the critical size in order to observe a motion on spatial scales of order $1/\sqrt{|\ln\varepsilon|}$ on which leapfrogging occurs. Due to the anisotropic nature of the problem leading to (1.1) in $x_{z}$-direction, one expects a less precise localization in $x_{z}$-direction which does not allow us to describe an entire crossing of vortex rings. Please notice that the radial confinement can be simply written as
$\big{|}r-X_{i,0,r}^{\varepsilon}\big{|}\leqslant\frac{1}{4}\min_{j\neq k}|X_{j,r}(0)-X_{k,r}(0)|$
by definition (1.3) of the $X_{i}(t)$.
The localization property and its link to the small-time assumption for our main result are discussed in view of possible future extensions more in detail in Remark 1.3 and a survey of previous results is provided in Section 1.3.
@@
Our method provides further information on the localization of vorticity. The center of potential vorticity of each vortex blob in the $2D$-reduction, namely the expression
(1.6) $b^{\varepsilon}_{i}(t)=\int_{{\mathbb{R}}^{2}_{+}}x\frac{x_{r}\omega^{\varepsilon}_{i}(t,x)}{\gamma}\mathrm{d}x$
can be localized as follows. We obtain in Proposition 5.4 as a consequence of the aforementioned localization properties that